# Finiteness of the set of associated primes for local cohomology modules of ideals via properties of almost factorial rings

### v2

@@ 1 Introduction
As mentioned above, regular rings of characteristic $p>0$, and regular semi-local rings containing a field are LC-finite. See [Mar], for more examples of an LC-finite ring.
@@
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
@@
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence of length at least $2$, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$, and suppose $\operatorname{depth}_{I}(R)=i-1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
@@
Over Noetherian local rings $R$ of small dimension, Marley[Mar] has shown several results regarding the finiteness of $\operatorname{Ass}H_{I}^{i}(M)$ for finite $R$-modules $M$. Using Marley’s result, which establishes the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, we deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ by applying the theorem above (see Proposition 5.3, Proposition 5.4).
@@ 2.1 Associated primes
Let $R$ be a ring, $M$ an $R$-module, and $P\in\operatorname{Ass}M$. Since there exists an injection $A/P\hookrightarrow M$, $0\to A_{P}/PA_{P}\to M_{P}$ is exact, where $A_{P}/PA_{P}\neq 0$. Thus, $M_{P}\neq 0$ and $P\in\operatorname{Supp}M$. This implies $\operatorname{Ass}M\subset\operatorname{Supp}M$.
@@ 3 Theoretical tools
In this section, we briefly summarize the Lewis’s approach to Question 1.2, which forms the basis for our discussion. We restate the question as follows: Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I\supset J$. Fix $i\geq 1$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite if $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite?
@@
Since a regular ring is locally almost factorial, Theorem 3.1 implies that $\operatorname{Ass}H_{I}^{2}(J)$ is always finite in the above question. Since $\operatorname{Ass}H_{I}^{1}(R/J)$ is finite, the question is true when $i=2$. The assertions for $i=3,4$ also hold under the hypothesis that $R/J$ is normal and locally almost factorial [Lew, Theorem 4.2(iii)]. The proof begins by reducing the problem to the case $\operatorname{depth}_{I}(R)\geq i-1$, using a generalized version of Hellus’s isomorphism (see [Lew, Theorem 4.1 and Corollary 4.3]). Furthermore,
@@
    In the case $\operatorname{depth}_{I}(R/J)=0$, the claim is true if $R/J$ is Cohen-Macaulay and a locally domain.
@@
Let $A$ be a Cohen-Macaulay ring, $I$ an ideal of $A$ with $\operatorname{depth}_{I}(A)=0$. If $A$ is a locally domain, $\sqrt{I}=\sqrt{aA}\cap I_{0}$ for some $a\in A$ and for some ideal $I_{0}$ with $\operatorname{depth}_{I_{0}}(A)>0$.
@@
Since $\operatorname{ht}I=\operatorname{depth}_{I}(A)=0$, we can write $\sqrt{I}=J\cap I_{0}$, in which $J$ is a finite intersection of minimal prime ideals of $A$, and $I_{0}$ is a finite intersection of height >0>0 prime ideals. Thus, $\operatorname{depth}_{I_{0}}(A)=\operatorname{ht}I_{0}>0$. To see that $J$ is a radical of some principal ideal of $A$, we show that every minimal prime ideal of $A$ is a radical of some principal ideal of $A$. Let $P_{1},\ldots,P_{t}$ be all the minimal prime ideals of $A$. Since $A$ is Noetherian and a locally domain,
@@
Thus, $e_{i}\notin P_{i}$, and $e_{j}\in P_{i}$ if $j\neq i$. In addition, $e_{1}+\cdots+e_{t}=1$, since this element corresponds to $(1,\ldots,1)$. Hence, $1-e_{i}\in P_{i}$, and it follows that $P_{i}$ is the only minimal element of $V((1-e_{i})A)$. Indeed, if not, there exists $P\in V((1-e_{i})A)$ such that $P\supset P_{j}$ for some $j\neq i$. Then, $1-e_{i},e_{i}\in P$, and $1\in P$, yielding a contradiction. ∎
@@
$R/J$ is Cohen-Macaulay, since $R$ is regular and $J$ is generated by an $R$-sequence. In addition, by the assumption that $R/J$ is normal, $R/J$ is a locally domain. For all of the cases above, the claim holds by Theorem 3.3. ∎
@@ 4 Main results
for some $x_{1},\ldots,x_{e}\in A$ such that $(x_{e}x_{1},x_{e}x_{2},\ldots,x_{e}x_{e-1})A\subset I_{0}$. Then, for any $A$-module $M$ and any $i>e$, $H_{I}^{i}(M)\simeq H_{x_{e}A\cap I_{0}}^{i}(M)$.
@@
Let $R$ be a Noetherian ring, $J$ an ideal generated by $R$-sequence of length at least $2$. Set $d=\operatorname{depth}_{I}(R)$, and let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that $\operatorname{Ass}H_{I}^{d}(R)$, $\operatorname{Ass}H_{I}^{d+1}(R)$ are finite, and that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then, for all $i\leq d+1$,
@@
is exact. If $i<d$, $H_{I}^{i}(R)=0$, and the claim immediately follows. If $i=d$, the claim follows from the assumption that $\operatorname{Ass}H_{I}^{d}(R)$ is finite. ∎
@@ 5 Applications
Let $(R,\mathfrak{m})$ be an $r$-dimensional regular local ring. $x_{1},\ldots,x_{r}\in\mathfrak{m}$ is a regular system of parameters if $\mathfrak{m}=(x_{1},\ldots,x_{r})$, or equivalently if the images of $x_{1},\ldots,x_{r}$ in $\mathfrak{m}/\mathfrak{m}^{2}$ are linearly independent over $R/\mathfrak{m}$. For any ideal $I$ of $R$, $R/I$ is regular if and only if $I$ is generated by a part of a regular system of parameters. Let $I\supset J$ be ideals of $R$ with $\operatorname{depth}_{J}(R)\geq 2$. When each of $I$ and $J$ is generated by a part of a regular system of parameters, we can take an $R/J$-sequence as stated in the claim of Theorem 4.3:
@@
It is true for $i\leq 2$, see Section 1. Fix $i\geq 3$. As we have seen in Section 3, we may assume $\operatorname{depth}_{I}(R)=i-1\geq 2$. Since $R/J$ is regular, for the cases $\operatorname{depth}_{I}(R/J)=0,1$, the claim holds by Theorem 3.3. Fix $e=\operatorname{depth}_{I}(R/J)\geq 2$. By Lemma 5.1, there are elements $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is regular. The claim also holds for the case $e\geq 2$, by Theorem 4.3. ∎
We apply the theorem to deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$, based on the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, which is guaranteed by Marley’s result[Mar]:
Proposition 5.3.
Proof.
@@
We may assume $i>0$, $\operatorname{ht}J\geq 2$, and $\dim R/J=\dim R-\operatorname{ht}J\leq 5$. If $\dim R/J\leq 4$, we can prove the claim using the same method as in Proposition 5.3. If $p\in J$ holds, $R/J$ is of equal characteristic; If $p\notin\mathfrak{m}^{2}+J$ holds, $R/J$ is unramified; in conclusion, $R/J$ is unramified. If $\dim R/J=5$, $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite for all $i>0$, by [Mar, Theorem 2.10]. For $i>0$, the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by Theorem 5.2. ∎
@@
We should emphasize in these propositions that it is not required for a regular local ring $R$ to be of equal characteristic. Indeed, if $R$ is of equal characteristic, the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by the Lyubeznik’s results (see [Lyu1] and [Lyu2]), without the assumptions on $R$ and $I$ in the above propositions :
@@ Acknowledgments
The author would like to express their gratitude to Professor Yasunari Nagai for clear guidance on how to proceed with this research and for many helpful discussions.

### v3

@@ 1 Introduction
As mentioned above, regular rings of characteristic $p>0$, and regular semi-local rings containing a field are LC-finite. See [BBLSZ] and [Lyu3], for more examples of an LC-finite ring.
@@
See [DM] and [Mar], for more examples of an $R$-module $M$ such that $\operatorname{Ass}H_{I}^{i}(M)$ is finite.
@@
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$. Does the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$ imply the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$?
@@
Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence of length at least $2$, and $I$ an ideal of $R$ containing $J$. Fix $i\geq 1$, and suppose $\operatorname{depth}_{I}(R)=i-1$. Does the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$ imply the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$?
@@
Over Noetherian local rings $R$ of small dimension, Marley[Mar] has shown several results regarding the finiteness of $\operatorname{Ass}H_{I}^{i}(M)$ for finite $R$-modules $M$. Using Marley’s result, which establishes the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, we deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ by applying the theorem above.
@@ 2.1 Associated primes
Let $R$ be a ring, $M$ an $R$-module, and $P\in\operatorname{Ass}M$. Since there exists an injection $A/P\hookrightarrow M$, we have an exact sequence $0\to A_{P}/PA_{P}\to M_{P}$, where $A_{P}/PA_{P}\neq 0$. Thus, $M_{P}\neq 0$ and $P\in\operatorname{Supp}M$. This implies $\operatorname{Ass}M\subset\operatorname{Supp}M$.
@@ 3 Theoretical tools
In this section, we briefly summarize the Lewis’s approach to Question 1.2, which forms the basis for our discussion. We restate the question as follows: Let $R$ be an LC-finite regular ring, $J$ an ideal of $R$ generated by an $R$-sequence, and $I\supset J$. Fix $i\geq 1$. Does the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$ imply the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$?
@@
Regarding the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ for locally almost factorial rings, we have the following result from [Lew].
@@
$\square$
Since a regular ring is locally almost factorial, Theorem 3.1 implies that $\operatorname{Ass}H_{I}^{2}(J)$ is always finite in the above question. Since $\operatorname{Ass}H_{I}^{1}(R/J)$ is finite, the question is true when $i=2$. The assertions for $i=3,4$ also hold under the hypothesis that $R/J$ is normal and locally almost factorial [Lew, Theorem 4.2(iii)]. The proof begins by reducing the problem to the case $\operatorname{depth}_{I}(R)\geq i-1$, using a generalized version of Hellus’s isomorphism (see [Lew, Theorem 4.1 and Corollary 4.3]). Furthermore, we use the following lemma from [Lew].
@@
    In the case $\operatorname{depth}_{I}(R/J)=0$, the claim is true if $R/J$ is Cohen-Macaulay and locally a domain.
@@
Let $A$ be a Cohen-Macaulay ring with $\dim A>0$, and $I$ an ideal of $A$ with $\operatorname{depth}_{I}(A)=0$. If $A$ is locally a domain, $\sqrt{I}=\sqrt{aA}\cap I_{0}$ for some $a\in A$ and for some ideal $I_{0}$ with $\operatorname{depth}_{I_{0}}(A)>0$.
@@
Since $\operatorname{ht}I=\operatorname{depth}_{I}(A)=0$, we can write $\sqrt{I}=J\cap I_{0}$, in which $J$ is a finite intersection of minimal prime ideals of $A$, and $I_{0}$ is a finite intersection of height >0>0 prime ideals. Thus, $\operatorname{depth}_{I_{0}}(A)=\operatorname{ht}I_{0}>0$. To see that $J$ is a radical of some principal ideal of $A$, we show that every minimal prime ideal of $A$ is a radical of some principal ideal of $A$. Let $P_{1},\ldots,P_{t}$ be all the minimal prime ideals of $A$. Since $A$ is Noetherian and locally a domain,
@@
Thus, $e_{i}\notin P_{i}$, and $e_{j}\in P_{i}$ if $j\neq i$. In addition, $e_{1}+\cdots+e_{t}=1$, since this element corresponds to $(1,\ldots,1)$. Hence, $1-e_{i}\in P_{i}$, and it follows that $P_{i}$ is the only minimal element of $V((1-e_{i})A)$. Indeed, if not, there exists $P\in V((1-e_{i})A)$ such that $P\supset P_{j}$ for some $j\neq i$. Then, $1-e_{i},e_{i}\in P$, and $1\in P$, yielding a contradiction. Since $J$ is a finite intersection of minimal prime ideals, we see that $J$ is of the form,
$J=\bigcap_{i}\sqrt{a_{i}A}=\sqrt{aA},\quad\text{where $a=\prod_{i}a_{i}$}.$
In the case where the depth equals $1$, we have the following result from [Lew].
@@
$\square$
The following result from [Lew] is useful when computing the local cohomology modules of ideals of a locally almost factorial ring.
@@
$R/J$ is Cohen-Macaulay, since $R$ is regular and $J$ is generated by an $R$-sequence. In addition, by the assumption that $R/J$ is normal, $R/J$ is locally a domain. For all of the cases above, the claim holds by Theorem 3.3. ∎
@@ 4 Main results
for some $x_{1},\ldots,x_{e}\in A$ such that $(x_{e}x_{1},x_{e}x_{2},\ldots,x_{e}x_{e-1})A\subset I_{0}$. Then, $H_{I}^{i}(M)\simeq H_{x_{e}A\cap I_{0}}^{i}(M)$ for any $A$-module $M$ and any $i>e$.
@@
Constructing an explicit example in which the quotient ring is locally almost factorial appears to be quite subtle. Indeed, even when $J$ is generated by an $R$-sequence, it seems difficult in general to determine whether the quotient $R/J$ is normal or factorial. At present, the main situation where the technical assumptions of Theorem 4.3 can be verified is the case of a regular local ring, as treated in Section 5.
@@
Let $R$ be a Noetherian ring, $J$ an ideal generated by an $R$-sequence of length at least $2$. Set $d=\operatorname{depth}_{I}(R)$, and let $I$ be an ideal containing $J$ such that $e=\operatorname{depth}_{I}(R/J)\geq 2$. Suppose that $\operatorname{Ass}H_{I}^{d+1}(R)$ are finite, and that there is an $R/J$-sequence $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is a locally almost factorial normal ring. Then, for all $i\leq d+1$,
@@
is exact. If $i<d$, then $H_{I}^{i}(R)=0$, and the claim immediately follows. If $i=d$, the claim follows from the fact that $\operatorname{Ass}H_{I}^{d}(R)$ is finite. Indeed, since $d=\operatorname{depth}_{I}(R)$, $H_{I}^{d}(R)$ is the first nonzero local cohomology modules of $R$ with respect to $I$. Then $\operatorname{Ass}H_{I}^{d}(R)$ is finite by [BrFa, proposition 2.1]. ∎
@@ 5 Applications
Let $(R,\mathfrak{m})$ be an $r$-dimensional regular local ring. $x_{1},\ldots,x_{r}\in\mathfrak{m}$ is a regular system of parameters if $\mathfrak{m}=(x_{1},\ldots,x_{r})$, or equivalently if the images of $x_{1},\ldots,x_{r}$ in $\mathfrak{m}/\mathfrak{m}^{2}$ are linearly independent over $R/\mathfrak{m}$. For any ideal $I$ of $R$, $R/I$ is regular if and only if $I$ is generated by a part of a regular system of parameters. Let $I\supset J$ be ideals of $R$ with $\operatorname{depth}_{J}(R)\geq 2$. When each of $I$ and $J$ is generated by a part of a regular system of parameters, we can take an $R/J$-sequence as stated in Theorem 4.3:
@@
It is true for $i\leq 2$, see Section 1. Fix $i\geq 3$. As we have seen in Section 3, we may assume $\operatorname{depth}_{I}(R)=i-1\geq 2$. Since $R/J$ is regular, for the cases $\operatorname{depth}_{I}(R/J)=0,1$, the claim holds by Theorem 3.3. Fix $e=\operatorname{depth}_{I}(R/J)\geq 2$. Then, there are elements $x_{1},\ldots,x_{e-1}\in I$ such that $R/(x_{1},\ldots,x_{e-1},J)$ is regular. The claim also holds for the case $e\geq 2$, by Theorem 4.3. ∎
Example 5.3.
Let $R=\mathbb{Z}_{p\mathbb{Z}}[[x_{1},\ldots,x_{n}]]$, a formal power series ring over a discrete valuation ring $\mathbb{Z}_{p\mathbb{Z}}$. $R$ is a regular local ring with its maximal ideal $\mathfrak{m}=(p,x_{1},\ldots,x_{n})$ and its residue field $\mathbb{Z}/p\mathbb{Z}$. A ring like $R$ is called an unramified ring of unequal characteristic (see below for the precise definition).
$R$ is known to be LC-finite (see [Lyu3] or [BBLSZ, Corollary 4.2]). Let $I\supset J$ be ideals, each of which is generated by a part of a regular system of parameters. Assume $p\in J$. Then $R/J$ is a regular local ring of characteristic $p$, and it is LC-finite. By Theorem 5.2, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any $i\geq 0$. Proposition 5.4 is a generalization of this example.
@@
Let $(R,\mathfrak{m})$ be an unramified regular local ring of unequal characteristic. Suppose that each of $I$ and $J$ is generated by a part of a regular system of parameters, and $p:=\operatorname{char}R/\mathfrak{m}\in J$. Then, $\operatorname{Ass}H_{I}^{i}(J)$ is finite for any $i\geq 0$.
Proof.
Since $R$ is LC-finite [Lyu3, Theorem 1] and $R/J$ is a regular local ring of characteristic $p$, the claim follows by Theorem 5.2. ∎
Next, we apply the theorem to deduce the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$, based on the finiteness of $\operatorname{Ass}H_{I}^{i-1}(R/J)$, which is guaranteed by Marley’s result[Mar]:
Proposition 5.5.
Proof.
Proposition 5.6.
@@
We may assume $i>0$, $\operatorname{ht}J\geq 2$, and $\dim R/J=\dim R-\operatorname{ht}J\leq 5$. If $\dim R/J\leq 4$, we can prove the claim using the same method as in Proposition 5.3. If $p\in J$ holds, $R/J$ is of equal characteristic; If $p\notin\mathfrak{m}^{2}+J$ holds, $R/J$ is unramified; in conclusion, $R/J$ is unramified. If $\dim R/J=5$, $\operatorname{Ass}H_{I}^{i-1}(R/J)$ is finite for all $i>0$, by [Mar, Theorem 2.10]. The finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by Theorem 5.2. ∎
@@
We should emphasize in these propositions that it is not required for a regular local ring $R$ to be of equal characteristic. Indeed, if $R$ is of equal characteristic, the finiteness of $\operatorname{Ass}H_{I}^{i}(J)$ follows by the Lyubeznik’s results (see [Lyu1] and [Lyu2]).
@@ Acknowledgments
The author is grateful to Professor Yasunari Nagai for valuable discussions and clear guidance throughout this research. The author also would like to thank the anonymous referees and the editor for their careful reading and helpful comments.